The annual salary for a certain job has a normal distribution with a mean of $54,00054,000 and a standard deviation of σ=$5000σ=$5000.
Answer #1

g tine we have to calculate Pl X L54215 which i denoted in Curve 2 Hire et) so y un 54 54000 50002 loo flon Z: 7 - 59000 unloS PLX>65) = L- PCX565) H 2 here N180, 12) Mio soplx 65)= |- P(XL 65 I NORM DIST (65, 80, 12, Extmutatives? - 1-0 = 1 = P(x>65